# Ahmed Hasanaath1, Aisha Alansari1, Ahmed Ashraf1, Chafik Salmane3, Hamzah Luqman1,2, Saad Ezzini1

### v1

@@ 4.3.1 Model Performance Across Task Categories
As outlined in Table 2, the DeepSeek family of models demonstrated strong capabilities across most Arabic NLP tasks, particularly excelling in SA (89.82%), StD (96.71%), and PD (94.14%) tasks. However, these models struggled most with PoS tagging, achieving a maximum F1-score of only 32.54% with 5-shot prompting, suggesting limitations in morphological analysis of Arabic text.
For NLG tasks, DeepSeek models showed competitive performance in MT (BLEU score of 50.27 with 5-shot prompting) but were outperformed by GPT-4o on PAR tasks. The ROUGE-L scores for SUM were relatively consistent across prompting strategies, indicating that this capability might be more inherent to the model’s pre-training rather than influenced by in-context examples.
@@ 4.3.3 Model Size and Architecture Effects
As demonstrated in Table 3, larger model sizes generally correlate with better performance. The largest DeepSeek R1 variant (671B) consistently outperforms smaller versions across most tasks, with an average performance improvement of 25.17 percentage points compared to the 14B variant in zero-shot setting. However, the relationship between model size and performance is not strictly linear, as evidenced in Table 2 by the QwQ-32B model underperforming by 10.62 percentage points compared to the DeepSeek R1-Q32B despite similar parameter counts.
@@ 6 Limitations
Despite the strengths of this study, several limitations should be acknowledged. First, the evaluation was limited to a subset of available datasets and target tasks, which may not fully capture the linguistic richness and variability of the Arabic language, including its diverse dialects, each with unique characteristics and challenges. Second, only a subset of LLMs was considered, potentially limiting the generalizability of the findings across different model architectures and training paradigms. Third, prompt selection was performed using a single LLM, which may introduce bias. Exploring the impact of prompts on other LLMs is important for a comprehensive evaluation of their performance. Finally, the prompting strategy was limited to three types of prompts. Although this was sufficient to establish baseline comparisons, it may not fully leverage the capabilities of each model. More advanced prompt engineering techniques, such as chain-of-thought prompting or opinion-based prompts, could lead to improved performance and deserve further exploration in future work.
A Data statistics
@@
Dataset Task(s) Train size Dev size Test size
AJGT SA 1,440 - 360
ArSarcasm DD, SarD 8,437 - 2,110
OSACT4 HSD, OSD 7,000 - 1,000
ANS-claim CD 3,185 987 456
ANS StD 2,652 756 379
NSURL-2019 PD 11,997 - 3,715
AraSum, EASC SUM 39,683 - 152
APPSD PAR 15,769 - 1,010
BOLT TRL 60,479 - 6,630
UNv1 MT 99,999 - 4,000
XTREME QA 14,805 - 921
Universal Dependencies v2.3 corpus POS 6,075 909 680
Arabic WSD Benchmark WSD 24,878 - 6,220
@@ B Prompt Strategies
In-Context Learning. For in-context learning, we evaluated models using both 3-shot and 5-shot configurations. Default model parameters from the API were used in this setting as well. To select the samples involved in the in-context learning experiments, we first estimated uncertainty scores using AraBERT’s [35] representations over the training data. The rationale behind using uncertain samples was to expose the model to more informative and challenging instances, which are more likely to enhance reasoning capabilities during inference [36, 37]. From the top 20 most uncertain samples, we manually selected 5 examples. These examples were then used for both 3-shot and 5-shot prompts. Manual curation ensured that examples exhibiting high uncertainty due to noise or ambiguity were excluded, focusing instead on those that were difficult yet coherent. From the top 20 most uncertain samples, we manually selected 5 examples. These examples were then used for both 3-shot and 5-shot prompts.
C Prompt Selection
@@
Instruction Selection. For each task, we designed a total of six prompt variants by crossing three styles with two languages (English and Arabic). The three prompt styles were: (1) Instructive, where the prompt uses an imperative tone (e.g., “Classify the sentiment of the following sentence.”); (2) Interrogative, where the prompt takes the form of a question (e.g., “What is the sentiment expressed in this sentence?”); and (3) Role-playing, which frames the instruction as if the model is performing a task in a role (e.g., “You are a sentiment analysis expert. Analyze the following sentence and determine its sentiment.”). Each of these three variants was written in both English and Arabic, resulting in six total prompts per task. To determine the most effective instruction style for each task, we evaluated performance using either the official dev set (if available) or a 300-example subset randomly sampled from the training set (see Appendix A Data statistics). The prompt with the highest score was selected for use in all downstream experiments, including both in-context learning and fine-tuning. See Figure 4 and Figure 5 for the complete list of the instructions we evaluated.
@@
Selecting Examples for Few-Shot Prompts. To select representative and informative examples for few-shot prompting, we adopted an uncertainty-based sampling strategy. Using AraBERT [35], we computed uncertainty scores across the training set of each task. Prior work has shown that exposing LLMs to more uncertain and challenging examples can improve performance by encouraging more robust reasoning [36, 37]. From the top 20 most uncertain samples, we manually selected 5 high-quality examples. These were used in both 3-shot and 5-shot prompt settings. Manual inspection allowed us to discard noisy or ambiguous examples, ensuring that the selected samples were difficult yet coherent and informative.
D Implementation Details
@@
For fine-tuning experiments, we focused on a subset of models from the DeepSeek family: R1-Q1.5B, R1-Q7B, and R1-Q14B. We fine-tuned the models using the Low-Rank Adaptation (LoRA) technique [38]. In our setup, LoRA was applied to an extended set of projection layers within the Transformer architecture, including the query, key, value, and output projections in the attention mechanism, as well as the gating and feedforward components (specifically, q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj). This broader application is motivated by findings from [39], which show that incorporating LoRA into MLP components improves adaptation performance while maintaining parameter efficiency. A LoRA rank of 4 was used. The maximum sequence length was set to 2048 tokens, with a batch size of 2 and gradient accumulation steps of 2, effectively simulating a batch size of 4. Fine-tuning was performed on an NVIDIA RTX A6000 GPU with 48 GB of VRAM. The training ran for up to 100,000 steps using the 8-bit AdamW optimizer with a cosine learning rate scheduler and a learning rate of 2e-4. Mixed-precision training was enabled via FP16 or BF16, depending on hardware support.
E Evaluation Results

### v4

@@ 4.3.1 Model Performance Across Task Categories
As outlined in Table 2, GPT o4-mini demonstrated strong capabilities across most Arabic NLP tasks, particularly excelling in SA (90.18%), DD (81.53%), HSD (71.14%), and PD (92.45%) tasks. However, this model struggled most with CD and StD, achieving a maximum F1-score of only 30.63% with zero-shot prompting for CD.
For NLG tasks, DeepSeek models showed competitive performance in MT (BLEU score of 26.39 with 5-shot prompting) but were outperformed by GPT-o4-mini on PAR tasks. The ROUGE-L scores for SUM were relatively consistent across prompting strategies, indicating that this capability might be more inherent to the model’s pre-training rather than influenced by in-context examples.
@@ 4.3.3 Model Size and Architecture Effects
As demonstrated in Table 3, larger model sizes generally correlate with better performance. The largest DeepSeek R1 variant (671B) consistently outperforms smaller versions across most tasks, with an average performance improvement of 25.17 percentage points compared to the 14B variant in the zero-shot setting. However, the relationship between model size and performance is not strictly linear, as evidenced in Table 2 by the QwQ-32B model underperforming by 10.62 percentage points compared to the DeepSeek R1-Q32B despite similar parameter counts.
@@ 6 Limitations
Despite the strengths of this study, several limitations should be acknowledged. First, the evaluation was limited to a subset of available datasets and target tasks, which may not fully capture the linguistic richness and variability of the Arabic language, including its diverse dialects, each with unique characteristics and challenges. Second, only a subset of LLMs was considered, potentially limiting the generalizability of the findings across different model architectures and training paradigms. Third, prompt selection was performed using a single LLM, which may introduce bias. Exploring the impact of prompts on other LLMs is important for a comprehensive evaluation of their performance. Finally, the prompting strategy was limited to three types of prompts. Although this was sufficient to establish baseline comparisons, it may not fully leverage the capabilities of each model. More advanced prompt engineering techniques, such as chain-of-thought prompting or opinion-based prompts, could lead to improved performance and deserve further exploration in future work. Furthermore, while we used a standardized prompt format across all models for consistency, we acknowledge that different models may be optimized for different prompting styles; and that prompt optimization per model is a potential direction for future work.
Acknowledgments
The authors would like to acknowledge the support received from the Saudi Data and AI Authority (SDAIA) and King Fahd University of Petroleum and Minerals (KFUPM) under the SDAIA-KFUPM Joint Research Center for Artificial Intelligence Grant JRC-AI-RFP-14.
Appendix A Data statistics
@@
Dataset Task(s) Train size Dev size Test size Full size
AJGT SA 1,440 - 360 1800
ArSarcasm DD, SarD 8,437 - 2,110 10,547
OSACT4 HSD, OSD 7,000 - 1,000 8,000
ANS-claim CD 3,185 987 456 4,628
ANS StD 2,652 756 379 3,787
NSURL-2019 PD 11,997 - 3,715 15,712
AraSum, EASC SUM 39,683 - 152 39,835
APPSD PAR 15,769 - 1,010 16,779
BOLT TRL 60,479 - 6,630 67,109
UNv1 MT 99,999 - 4,000 103,999
XTREME QA 14,805 - 921 15,726
Universal Dependencies v2.3 corpus POS 6,075 909 680 7,664
Arabic WSD Benchmark WSD 24,878 - 6,220 31,098
@@ Appendix B Prompt Strategies
In-Context Learning. For in-context learning, we evaluated models using both 3-shot and 5-shot configurations. Default model parameters from the API were used in this setting as well. To select the samples involved in the in-context learning experiments, we first estimated uncertainty scores using AraBERT’s antoun2020arabert representations over the training data. The rationale behind using uncertain samples was to expose the model to more informative and challenging instances, which are more likely to enhance reasoning capabilities during inference nguyen2022measure; zhu2009active. From the top 20 most uncertain samples, we manually selected 5 examples. These examples were then used for both 3-shot and 5-shot prompts. Manual curation ensured that examples exhibiting high uncertainty due to noise or ambiguity were excluded, focusing instead on those that were difficult yet coherent. From the top 20 most uncertain samples, we manually selected 5 examples. These examples were then used for both 3-shot and 5-shot prompts.
Appendix C Prompt Selection
@@
Instruction Selection. For each task, we designed a total of six prompt variants by crossing three styles with two languages (English and Arabic). The three prompt styles were: (1) Instructive, where the prompt uses an imperative tone (e.g., “Classify the sentiment of the following sentence.”); (2) Interrogative, where the prompt takes the form of a question (e.g., “What is the sentiment expressed in this sentence?”); and (3) Role-playing, which frames the instruction as if the model is performing a task in a role (e.g., “You are a sentiment analysis expert. Analyze the following sentence and determine its sentiment.”). Each of these three variants was written in both English and Arabic, resulting in six total prompts per task. To determine the most effective instruction style for each task, we evaluated performance using either the official dev set (if available) or a 300-example subset randomly sampled from the training set (see Appendix A). The prompt with the highest score was selected for use in all downstream experiments, including both in-context learning and fine-tuning. See Figure 4 and Figure 5 for the complete list of the instructions we evaluated.
@@
Selecting Examples for Few-Shot Prompts. To select representative and informative examples for few-shot prompting, we adopted an uncertainty-based sampling strategy. Using AraBERT antoun2020arabert, we computed uncertainty scores across the training set of each task. Prior work has shown that exposing LLMs to more uncertain and challenging examples can improve performance by encouraging more robust reasoning nguyen2022measure; zhu2009active. From the top 20 most uncertain samples, we manually selected 5 high-quality examples. These were used in both 3-shot and 5-shot prompt settings. Manual inspection allowed us to discard noisy or ambiguous examples, ensuring that the selected samples were difficult yet coherent and informative.
Appendix D Implementation Details
@@
For fine-tuning experiments, we focused on a subset of models from the DeepSeek family: R1-Q1.5B, R1-Q7B, and R1-Q14B. We fine-tuned the models using the Low-Rank Adaptation (LoRA) technique hu2022lora. In our setup, LoRA was applied to an extended set of projection layers within the Transformer architecture, including the query, key, value, and output projections in the attention mechanism, as well as the gating and feedforward components (specifically, q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj). This broader application is motivated by findings from hayou2024lora_plus, which show that incorporating LoRA into MLP components improves adaptation performance while maintaining parameter efficiency. A LoRA rank of 4 was used. The maximum sequence length was set to 2048 tokens, with a batch size of 2 and gradient accumulation steps of 2, effectively simulating a batch size of 4. Fine-tuning was performed on an NVIDIA RTX A6000 GPU with 48 GB of VRAM. The training ran for up to 100,000 steps using the 8-bit AdamW optimizer with a cosine learning rate scheduler and a learning rate of 2e-4. Mixed-precision training was enabled via FP16 or BF16, depending on hardware support.
Appendix E Evaluation Results